OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (2, 5) and (−1, 8)? y + 5 = 1(x + 2) y + 8 = −1(x − 1) y − 5 = −1(x − 2) y − 8 = 1(x − 2)

OpenStudy (anonymous):

I am clueless on this one lol....

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

point slope form :- y-y1 = m(x-x1)

ganeshie8 (ganeshie8):

you're given two points (2, 5) and (−1, 8)

ganeshie8 (ganeshie8):

first find slope

OpenStudy (anonymous):

sorry was afk

ganeshie8 (ganeshie8):

okie

OpenStudy (anonymous):

(y2 - y1) / (x2 - x1) (8 - 5) / (-1 - 2) 3 / -3

ganeshie8 (ganeshie8):

yea can it be simplified further ? :)

OpenStudy (anonymous):

3/-3... -1

OpenStudy (anonymous):

slope of -1

ganeshie8 (ganeshie8):

point slope form :- y-y1 = m(x-x1)

ganeshie8 (ganeshie8):

m = -1

ganeshie8 (ganeshie8):

pick a point :- (2, 5) x1 y1

ganeshie8 (ganeshie8):

plug them above

OpenStudy (anonymous):

y − 5 = −1(x − 2)

ganeshie8 (ganeshie8):

Yup !

OpenStudy (anonymous):

:D